OpenStudy (anonymous):

5. Find the 8th term in the geometric sequence with t4 = 48 and t6 = 192. a. + or - 1536 b.768 c. + or - 384 d. 192

OpenStudy (anonymous):

b. 768

OpenStudy (anonymous):

egjactly t1=6 t2=12 t3=24 t4=48 t5=96 t6=196 so each step the previous term is multiplied by 2 then t8=t6*4 or 192*4=768

OpenStudy (anonymous):

@theyatin Maybe 2 or -2 ...

OpenStudy (anonymous):

nope only 2

OpenStudy (anonymous):

take -2 ...it works also !

OpenStudy (anonymous):

it doesnt. it will make the multiple negative. . . dont consider it. . .

OpenStudy (anonymous):

Sorry ! But you have problem with negative numbers ...?! t7=192*2 or 192*-2 !

OpenStudy (anonymous):

you made a mistake ...Sorry !

OpenStudy (anonymous):

let it be neemo typing mistake. . . 96*2=192 96*-2=-193 there is a huge difference if u consider -2

OpenStudy (anonymous):

oh my god calm down please : ))))))))))))

OpenStudy (anonymous):

2and -2 give both the same result for t8 t10 t12 but for t7 t9 t10 you can't really know !!! sorry, Summer1nthex1 ...... I ' m sure ...you can ask your teacher ; if you want :D :D